3· value: 2.27 points Media Bias Inc. issued bonds 10 years ago at $1,000 per bond. These bonds had a 30- year life when issuCan you show me steps on how to solve this, please? Thank you!
Answer #1

Total return = real rate of return + inflation premium + risk premium

Total return = 4% + 5% + 2% = 11%

Coupon = 0.13 * 1,000 = 130

New price of bond = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

New price of bond = 130 * [1 - 1 / (1 + 0.11)20] / 0.11 + 1,000 / (1 + 0.11)20

New price of bond = 130 * 7.963328 + 124.033907

New price of bond = $1,159.27

Keys to use in a financial calculator: FV 1000, PMT 130, N 20, I/Y 11, CPT PV